CALIFORNIA STATE UNIVERSITY, HAYWARD DEPARTMENT OF STATISTICS

Statistics 3401 Introduction to Probability Theory I Fall 2001

Lecture: MW 6-8, NSc 207
Instructor: Prof. Eric A. Suess  Office Hours: MW 5-6pm

NSc 319 510-885-3879 esuess@csuhayward.edu www.telecom.csuhayward.edu/∼esuess/

Required Text:

• Wackerly, Mendenhall, and Scheaffer: Mathematical Statistics with Applications, 6th ed., Duxbury, 2002.

Topics:

In this course we will cover the fundamentals of Probability Theory.

• Rules of Probability

• Counting Rules

• Discrete and Continuous probability distributions

• Expectation and Variance of random variables

• Moment Generating Functions

• Multivariate probability distributions

• Marginal and Conditional probability distributions

• Functions of random variables

• Sampling Distributions

• Central Limit Theorem

Homework:

Weekly homework assignments will be assigned and collected on Mondays. The assignments will include problems from the textbook and will sometimes require the use of computer software.

Computers:

The Windows software, R, will be introduced and used for some of the homework assignments. http://www.r-project.org/

References:

• Venerables and Smith: Notes on R: A Programming Environment for Data Analysis and Graphics

Grading:

Homework 40%, Midterm I 20%, Midterm II 20%, Final 20%